The volume of a cone is 360cm3. If the height of the cone is 30cm,find the length of the base
1
Expert's answer
2012-08-30T11:20:10-0400
Assume that the base is a circle of radius r, so the area of the base is
S
= pi * r^2

The height of the cone is h=30 cm.

Then the volume is
equal to
V = h * S / 3 = 360

pi * r^2 * h = 3
V

whence

r = square_root ( 3 V/(pi*h) )

Therefore
the length of the base is

L = 2 * pi * r =
= 2 * pi *
square_root ( 3 V/(pi*h) )
= 2 * square_root ( 3 V pi /h )

= 2 * square_root ( 3 * 360 * pi / 30 )
= 2 * square_root ( 36 pi )

= 12 * square_root ( pi )
= 21.2694 cm